# Improving Equivariant Model Training via Constraint Relaxation

### v1

@@ Abstract
Equivariant neural networks have been widely used in a variety of applications due to their ability to generalize well in tasks where the underlying data symmetries are known. Despite their successes, such networks can be difficult to optimize and require careful hyperparameter tuning to train successfully. In this work, we propose a novel framework for improving the optimization of such models by relaxing the hard equivariance constraint during training: We relax the equivariance constraint of the network’s intermediate layers by introducing an additional non-equivariance term that we progressively constrain until we arrive at an equivariant solution. By controlling the magnitude of the activation of the additional relaxation term, we allow the model to optimize over a larger hypothesis space containing approximate equivariant networks and converge back to an equivariant solution at the end of training. We provide experimental results on different state-of-the-art network architectures, demonstrating how this training framework can result in equivariant models with improved generalization performance.
@@ 1 Introduction
  • We present a novel training framework that can improve the performance of equivariant neural networks by relaxing the equivariance constraint during training and projecting back to the space of equivariant models during testing.
@@ 3 Method
In this work we focus on a family of models as described above—that are defined through a composition of simpler equivariant linear layers. During standard training the linear layers are optimized over the set of intertwiners $\mathcal{H}_{i}$, i.e. the set of linear maps between the representations $(V_{i},\rho_{V_{i}})$ and $(V_{i+1},\rho_{V_{i+1}})$ that have the equivariance property as stated in Equation 1. The set of intertwiners is only a subset of the set of all possible linear maps from $V_{i}$ to $V_{i+1}$, and as a result, they have a reduced number of free parameters. We propose to facilitate training over this constrained space by relaxing the constraint imposed on the intermediate layers and optimizing over a larger hypothesis space $\tilde{H}$ which is a superset of the set of equivariant models $H\subset\tilde{H}$. A trivial approach is to completely relax the constraint and solely optimize over the larger set containing all models. The problem with such an approach is that it completely abandons the concept of equivariance and all the attendant generalization benefits. Consequently, to expand the hypothesis space while keeping the benefits of equivariant models, we need a relaxation such that:
@@ 4.1 Equivariant Point Cloud Classification
To apply our method we relax the Vector Neurons linear layer by following the methodology described in Section 3.4. For both networks we set the regularization term $\lambda_{reg}=0.01$, which is a value we use in all of the following experiments. We provide a more detailed description of the training parameters in Appendix A.1. Furthermore, in Appendix A.2 we describes the process of choosing the hyperparameter $\lambda_{reg}$ and show the method’s robustness to its value. Figure 2 showcases how applying our proposed framework benefits the training of both networks. Specifically, for the case of the smaller and less performant PointNet, we can see an even larger performance increase over the baseline. These results show how the performance benefits of our optimization framework increase in smaller under-parametrized networks, an effect we investigate further in Section 4.2. In Appendix A.3 we provide additional details on the computational and memory overhead of our proposed optimization, showcasing that while additional parameters are introduced during training the overhead in the training time is limited.
@@ A.1 Training Details
  • Point Cloud Classification: We use as baselines the VN-PointNet and VN-DGCNN network architectures described in the work of Deng et al. (2021). For the relaxed version of VN-PoitNet we train for 250 epochs using the Adam optimizer (Kingma & Ba, 2015), with an initial learning rate of $10^{-3}$, that we decrease every 20 epochs by a factor of 0.7, and weight decay equal to $10^{-4}$. For the relaxed version of VN-DGCNN we train for 250 epochs using stochastic gradient descent, with an initial learning rate of $10^{-3}$, that we decrease using cosine annealing, and weight decay equal to $10^{-4}$. The batch size used was 32.
@@ A.4 Ablation on $\theta$ Scheduling and Equivariant projection
A.5 Limitations

### v2

@@ Abstract
Equivariant neural networks have been widely used in a variety of applications due to their ability to generalize well in tasks where the underlying data symmetries are known. Despite their successes, such networks can be difficult to optimize and require careful hyperparameter tuning to train successfully. In this work, we propose a novel framework for improving the optimization of such models by relaxing the hard equivariance constraint during training: We relax the equivariance constraint of the network’s intermediate layers by introducing an additional non-equivariant term that we progressively constrain until we arrive at an equivariant solution. By controlling the magnitude of the activation of the additional relaxation term, we allow the model to optimize over a larger hypothesis space containing approximate equivariant networks and converge back to an equivariant solution at the end of training. We provide experimental results on different state-of-the-art network architectures, demonstrating how this training framework can result in equivariant models with improved generalization performance. Our code is available at https://github.com/StefanosPert/Equivariant_Optimization_CR
@@ 1 Introduction
  • We present a novel training framework that can improve the performance of equivariant neural networks by relaxing the equivariance constraint during training and projecting back to the space of equivariant models during testing (as shown in Figure 1).
@@ 3 Method
In this work we focus on a family of models as described above—that are defined through a composition of simpler equivariant linear layers. During standard training the linear layers are optimized over the set of intertwiners $H_{i}$, i.e. the set of linear maps between the representations $(V_{i},\rho_{V_{i}})$ and $(V_{i+1},\rho_{V_{i+1}})$ that have the equivariance property as stated in Equation 1. The set of intertwiners is only a subset of the set of all possible linear maps from $V_{i}$ to $V_{i+1}$, and as a result, they have a reduced number of free parameters. We propose to facilitate training over this constrained space by relaxing the constraint imposed on the intermediate layers and optimizing over a larger hypothesis space $\tilde{H}$ which is a superset of the set of equivariant models $H\subset\tilde{H}$. A trivial approach is to completely relax the constraint and solely optimize over the larger set containing all models. The problem with such an approach is that it completely abandons the concept of equivariance and all the attendant generalization benefits. Consequently, to expand the hypothesis space while keeping the benefits of equivariant models, we need a relaxation such that:
@@ 4.1 Equivariant Point Cloud Classification
To apply our method we relax the Vector Neurons linear layer by following the methodology described in Section 3.4. For both networks we set the regularization term $\lambda_{reg}=0.01$, which is a value we use in all of the following experiments. We provide a more detailed description of the training parameters in Appendix A.1. Furthermore, in Appendix A.2 we describe the process of choosing the hyperparameter $\lambda_{reg}$ and show the method’s robustness to its value. Figure 2 showcases how applying our proposed framework benefits the training of both networks. Specifically, for the case of the smaller and less performant PointNet, we can see an even larger performance increase over the baseline. These results show how the performance benefits of our optimization framework increase in smaller under-parametrized networks, an effect we investigate further in Section 4.2. In Appendix A.3 we provide additional details on the computational and memory overhead of our proposed optimization, showcasing that while additional parameters are introduced during training the overhead in the training time is limited.
@@ A.1 Training Details
  • Point Cloud Classification: We use as baselines the VN-PointNet and VN-DGCNN network architectures described in the work of Deng et al. (2021). For the relaxed version of VN-PoitNet we train for 250 epochs using the Adam optimizer (Kingma & Ba, 2015), with an initial learning rate of $10^{-3}$, that we decrease every 20 epochs by a factor of 0.7, and weight decay equal to $10^{-4}$. For the relaxed version of VN-DGCNN we train for 250 epochs using stochastic gradient descent, with an initial learning rate of $10^{-1}$, that we decrease using cosine annealing, and weight decay equal to $10^{-4}$. The batch size used was 32.
@@ A.4 Ablation on $\theta$ Scheduling and Equivariant projection
A.5 Additional comparison with Methods using Equivariant Adaptation/Fine Tuning
As described in Section 2, while our work focuses on improving the optimization of equivariant networks, the works of Mondal et al. (2023) (Equiv-Adapt) and Basu et al. (2023b) (Equi-Tuning) focus on circumventing the need of optimizing equivariant network by performing equivariant adaptation or fine tunining of non-equivariant models. Since such methods have a different focus, mainly utilizing already pre-trained non-equivariant models to solve equivariant tasks, a straightforward comparison can be challenging. Nevertheless, in Table 4 we provide a comparison with our proposed method on the task of sparse point cloud classification. We can see that our method outperforms Equiv-Adapt and has performance close to the one achieved by Equi-Tuning which requires multiple forward passes during inference.
Table 4: Comparison of our proposed method with previous works performing equivariant adaption or finetuning, on ModelNet40 classification (Base model: VN-PointNet). Here it is important to note that in the case of Equi-Tuning, equivariance is achieved by group averaging. As a result, during inference the model is required to perform multiple forward passes, which slows down the method’s inference.
Equiv-Adapt Equi-Tuning Original VNN Ours
66.3% 74.9% 66.4% 74.5%
A.6 Limitations